OpenStudy (styxer):

If a function f(x) in the domain x E [0,2] is f(x) = |x-1| + |x²-2x|, then the minimum value is (?) and the maximum value is (??)

OpenStudy (eliesaab):

consider two cases \[ 0\le x \le 1\\ 1<x\leq 2 \]

OpenStudy (eliesaab):

In the first case \[ f(x)= 1-x+ 2 x-x^2=-x^2 +x +1 \]

OpenStudy (eliesaab):

For the second case \[ f(x) = x-1 +2x-x^2=-x^2+3x-1 \]

OpenStudy (eliesaab):

Why is that and how can you finish the problem?

OpenStudy (styxer):

@eliesaab I can use the Xv and Yv formulas, but what about the "x" range? Doesn't it interfer? Or I can make sure about the range by calculating the Xv?

OpenStudy (styxer):

@eliesaab the higher Yv value in the two equations will be the maximum, and the other is the minimum?

OpenStudy (loser66):

To me, you need put everything in logic, like this for |x -1| : a) = x-1 if x-1 >0, or x >1 b) = 1-x if x-1 <0 , or x <1

OpenStudy (loser66):

for |x^2-2x| c) = x^2 -2x if x^2 -2x = x(x-2) >0 There are 2 cases here both x > 0 and x -2 >0 ( x>2) makes the product >0 Or both x <0 and x -2 <0 , this makes the product > 0 also. However, we reject this case because it is out of the domain. d) -(x^2-2x) =2x -x^2 if x(x-2) <0 For this case, we need x >0 and x -2 <0 , that is 0 < x<2 . We need this case. for x < 0 and x-2 >0 , reject.

OpenStudy (loser66):

Now, combine!! [0, 1] case b + case d, then we have 1-x - x^2+ 2x = -x^2 +x -1

OpenStudy (loser66):

[1,2] we have case a + case d) that i x-1 +2x -x^2= -x^2+3x-1

OpenStudy (loser66):

For both interval, take derivative to find the max, min Then, find f(0), f(1) f(2) to find the min.

OpenStudy (tkhunny):

You may wish to consider the equivalent form, f(x) = |x-1|+x|x-2|. This rather clearly marks where you should look. Isn't it a joy when things are on the boundary and calculus doesn't help?

OpenStudy (styxer):

@loser66 so, both equations gives 5/4 as max. value and f(0),f(1),f(2) gives 1 (it's the minimum,right?) , we have to calculate for 0,1 and 2 because 1 is root for the 1st module and 0 and 2 are roots for the 2nd module? I didn't quite understand how you got 4 values on the |x²-2x| part... Can you explain me, please? @tkhunny that's why the minimum values are on (for x) 0,1 and 2?

OpenStudy (eliesaab):

Look at the graph

OpenStudy (eliesaab):

Three minimum at x=0,1,and 2. That minimum is 1 at x=.5 and 1.5 we have a maximum of 1.25

OpenStudy (styxer):

@eliesaab I see now. Thanks guys/girls!